Circular No. 8835

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 WD_{4} (LEMMON) E. Guido and G. Sostero, Castellammare di Stabia, Italy, report that

their remote CCD observations with a 0.15-m f/6 reflector at the ‘Rent-a-Scope’ Observatory at Biggera Waters, Queensland, show the Mt. Lemmon discovery 2006 WD_{4} (cf. MPEC 2006-W61, MPC 189005; discovery observation tabulated below) to be a comet. The co-addition of eleven unfiltered 60-s exposures on Apr. 30.80 UT revealed the presence of a spiral-like symmetric coma, spanning almost 20′′, from p.a. 240^{o} to p.a. 30^{o} (in a counterclockwise direction); the total magnitude was close to 14.2, and that of the condensation was 15.3–15.4. The coaddition of nine similar exposures (but in less satisfactory conditions) on May 2.80 showed the presence of an elliptical coma measuring nearly 12′′ × 18′′, the major axis tilted along a line close to p.a. 235^{o}; the total mag was about 14.1 and that of the condensation 15.4.

2006 UT  α_{2000}  δ_{2000}  Mag.
Nov. 20.32650  3^{h}32^{m}32.^{s}53  +28^{o}02′22′′.0  20.3

The following orbital elements, together with the Apr. 30–May 2 astrometry, full residuals, and an ephemeris, are given on MPEC 2007-J12.

Epoch = 2007 Apr. 10.0 TT

T = 2007 Apr. 28.3951 TT  ω = 292.^{o}6928
e = 0.998994  Ω = 226.7908 }2000.0
q = 0.591207 AU  i = 152.7038

With reference to observations by F. Bernardi with the 2.24-m University of Hawaii reflector on 2006 Dec. 25.4 UT, D. J. Tholen wrote that, although the object appeared stellar to the eye, the CCD trail of 2006 WD_{4} was wider than the trails of two dozen stars nearby. In the second of the 600-s exposures, the mean stellar trail width was 2.08 ± 0.05 pixels (0′′.91); that of 2006 WD_{4} (FWHM) was 2.3 pixels (1′′.01), about 0.1 pixel wider than the widest of the stellar sources. The results were considered “sufficiently suspicious to warrant further observations as the object approaches the sun”.

COMET 187P/LINEAR Comet P/1999 J5 = 2007 E3 (cf. IAUC 8827) has been assigned the

permanent number 187P (cf. MPC 59555).

2007 May 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green